OpenStudy (anonymous):

Simplify each expression. Assume that all variables are positive. 1) \[(4x ^{-2}y^4)^{-1/2}\] 2) \[(8ab^2)^{-1/2}(8ab^2)^{1/2}\] 3) \[(s ^{2/5}t ^{1/3})(s ^{1/2}t ^{1/2})\] I really have no idea how to do them. Explain and show steps please!

OpenStudy (callisto):

1. (4x^-2 y^4 )^-1/2 = 1/ [4x^-2 y^-4) ^ (1/2) ] = (x^2)^(1/2) / [4y^-4) ^ (1/2) ] = x/(2y^2)

OpenStudy (callisto):

2. (8ab^2)^(-1/2) (8ab^2)^(1/2) = (8ab^2)^(1/2) / (8ab^2)^(1/2) =1

OpenStudy (callisto):

well you need to know that x^(-a) = 1/ (x^a)

OpenStudy (callisto):

3. 8^(2/5) t^(1/3) 8^(1/2) t^(1/2) = 8^ (2/5 + 1/2 ) t^(1/3 + 1/2) = 2^(3x9/10) t^(5/6) = 2^(27/10) t^(5/6)

OpenStudy (anonymous):

The answer for 2 is just 1?

OpenStudy (callisto):

yes =(8ab^2)^(-1/2 + 1/2) = (8ab^2)^(0) =1 note that a^0 = 1

OpenStudy (anonymous):

or @campbell_st @amistre64

OpenStudy (campbell_st):

when multiplying the rule is add the powers of the same pronumeral for s its 2/5 + 1/2 = 9/10 for t its 1/3 + 1/2 = 5/6 \[s^{9/10}t^{5/6}\]

OpenStudy (anonymous):

so that's all?

OpenStudy (campbell_st):

yep... index laws... the rule applies for multiplication

OpenStudy (anonymous):

@campbell_st are all the other answer good too?

OpenStudy (campbell_st):

I haven't looked at them... been busy... sorry... I don't have time

OpenStudy (anonymous):

@campbell_st it's okay look at them when you have time and let me know please